The width of a rectangle is 4 feet less than the length. The area is 320ft^2. Find the width of the rectangle.
___feet
Thank you

Answer by Nate(3500)

Width(w) = l - 4
Length(l) = l
Length * Width = Area
(l)(l - 4) = 320
l^2 - 4l - 320 = 0
(l - 20)(l + 16) = 0
l+=+20 or l+=+-16
Since your length can not be negative, your length is: l+=+20
Plug:
Width = l - 4
Width = 20 - 4
Width = 16